OpenStudy (anonymous):

Solve 2(x + 1) = 2x + 5. All real numbers No solution 0 3

OpenStudy (anonymous):

@ParthKohli please help

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

can you simplify the left hand side first?

OpenStudy (anonymous):

No, I don't believe so

OpenStudy (anonymous):

2(x + 1), you need to distribute here, are you familiar with distributing?

OpenStudy (anonymous):

actually yes.

OpenStudy (anonymous):

ok so please distribute to simplify that

OpenStudy (anonymous):

2x + 2 right ?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

so now you have 2x + 2 = 2x + 5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now subtract 2x from both sides

OpenStudy (anonymous):

what do you end up with?

OpenStudy (anonymous):

ummm... I think it is 2x = 2x + 3

OpenStudy (anonymous):

I think I did it wrong

OpenStudy (anonymous):

well that is correct, if you subtract 2 from both sides, but i was saying to subtract 2x from both sides

OpenStudy (anonymous):

so if you have 2x + 2 = 2x + 5, and you subtract 2x from both sides you get 2 = 5, since 2 does not equal 5 this does not have any solutions

OpenStudy (michele_laino):

hint: if we apply the distributive property of multiplication over addition at left side, we can write this: 2x+2=2x+5

OpenStudy (anonymous):

does this make sense?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

perfect

OpenStudy (anonymous):

Thanks guys for all your help!

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

I need help with another one, can you guys help?

OpenStudy (anonymous):

sure, close this one and post up another